OpenStudy (anonymous):

The length of a rectangle is 4 inches more than its width, and its perimeter is 34 inches. Find the length and width of the rectangle. If w = the width of the rectangle, then the length equals _____. w - 4 w + 4 4w

OpenStudy (anonymous):

@ANIMALOVER02

OpenStudy (just_one_last_goodbye):

We suppose that the width of rectangle is x. Then the length would be (x+4). The question says that the perimeter is 34 inches. Perimeter is the total length of all four sides of a rectangle. So: x+(x+4)+x+(x+4)=34 4x+16=34 4x=34-16 x=18/4 x=4.5 inches

OpenStudy (just_one_last_goodbye):

Oops. I made a mistake in the calculations: This would be; 4x+8=34 4x=34-8 x=26/4 x=6.5 inches

OpenStudy (just_one_last_goodbye):

So since the width is 6.5 inches. The length would be (6.5+4)= 10.5 inches. Sorry for that mistake.

OpenStudy (anonymous):

so what is it?

OpenStudy (just_one_last_goodbye):

I think it's w + 4

OpenStudy (just_one_last_goodbye):

Not sure tho